OpenStudy (anonymous):

what is the greatest common factor of 15a^4 and 14a^2?

OpenStudy (unklerhaukus):

If you fully factorise those terms \[15a^4 = 3\cdot5\cdot a\cdot a\cdot a\cdot a\\ 14a^2= 2\cdot 7\cdot a\cdot a\] Which factors are common?